MIDC/Data Quality Statements
Total Records: 40, displaying 1-25
[Next Records]
R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
1BMSPhotometer TWC05/05/202119:00:00TWC Photometer is down due to an unknown problem. The cables were troubleshot on 6/24 and no problem was found. Desiccant was checked and no moisture was found. After reassembly TWC was working fine, problem is intermittent.
2BMSPhotometer TWC03/23/202100:00:0003/24/202112:46:00TWC Photometer is down due to a problem with signal cable between instrument and outdoor junction box. The cables were troubleshot on 3/24 and no problem was found. After reassembly TWC was working fine, problem could be intermittent.
3BMSPhotometer TWC11/17/202007:21:0011/17/202007:45:00Due to a scheduled NREL/SRRL power outage, to replace a surge protection device on the SRRL building, the following BMS instruments were down: 501A-UVA, TSP-700, UVSAT, UVSBT, SolarSIM-D2, CUVA2, WISERs (Horiz;2-Axis;1-Axis), CUVB2, UVB-1, 501A-UVB, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, ASI-16.
4BMSPhotometer TWC03/09/201908:40:0003/09/201912:59:00Due to a scheduled NREL/SRRL power outage, to install power meters on the SRRL building, the following BMS instruments were down: 501A-UVA, TSP-700, UVSAT, UVSBT, SolarSIM-D2, CUVA2, WISERs, CUVB2, UVB-1, 501A-UVB, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, ASI-16.
5BMSPhotometer TWC03/09/201908:40:0003/09/201913:01:00Due to a scheduled NREL/SRRL power outage, to install power meters on the SRRL building, The following BMS instruments were not pointing at the sun during this time: Direct NIP, LI-201, LI-190, NIP RG780, TWC, DR02; and the following instruments were un-shaded: CG4.
6BMSPhotometer TWC05/11/201823:56:0005/14/201807:05:00A GFCI on the North tracker post was tripped during rain shower. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC.
7BMSPhotometer TWC01/17/201707:15:0001/17/201708:56:00Due to a tripped GFCI powering the Sci-Tec tracker, The following BMS instruments were not pointing at the sun during this time: Direct CHP1 #2, NIP, DR02, LI-201, LI-190, NIP RG780, TWC; and the following instruments were un-shaded: Diffuse CM22 #1, CMP11, CG4.
8BMSPhotometer TWC05/02/201509:00:0005/02/201518:53:00Due to a snagged signal cable on the Sci-Tec tracker, The following BMS instruments were not pointing at the sun during this time: Direct NIP #1, LI-201, LI-190, NIP RG780, TWC; and the following instruments were un-shaded: Diffuse CM22, CMP11, CG4.
9BMSPhotometer TWC04/25/201507:20:0004/25/201509:39:00Due to a scheduled NREL/SRRL power outage, to remove wiring from the building ATS, the following BMS instruments were down: TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100, MS-7XX, LI-1800.
10BMSPhotometer TWC05/30/201416:18:0005/31/201411:30:00A GFCI on the North tracker post was tripped due to water logged ventilators (caused by rain showers). Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
11BMSPhotometer TWC12/28/201306:46:0012/28/201315:59:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
12BMSPhotometer TWC12/21/201307:00:0012/21/201316:22:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
13BMSPhotometer TWC12/21/201307:00:0012/21/201316:22:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, The following BMS instruments were not pointing at the sun during this time: Direct NIP #1, LI-201, LI-190, NIP RG780, TWC; and the following instruments were un-shaded: Diffuse CM22, CUV4, CG4. The Middle Sci-Tec tracker for these instruments were temporarily put on a secondary UPS which failed immediately.
14BMSPhotometer TWC09/13/201316:08:0009/14/201309:55:00A GFCI on the North tracker post was tripped due to water logged ventilators (caused by rain showers). Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
15BMSPhotometer TWC07/13/201316:23:0007/14/201308:29:00A GFCI on the North tracker post was tripped due to water logged ventilators (caused by rain showers). Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
16BMSPhotometer TWC05/01/201313:27:0005/02/201306:40:00A GFCI on the North tracker post was tripped due to thunderstorms. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
17BMSPhotometer TWC03/02/201313:06:0003/02/201318:24:00Due to a scheduled NREL power outage, to upgrade campus power capacity, The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4. The Sci-Tec tracker for these insruments are protected by UPS, however the battery depleted at 13:06 MST (the power outage started at 8:06 MST)
18BMSPhotometer TWC03/02/201308:06:0003/02/201318:24:00Due to a scheduled NREL power outage, to upgrade campus power capacity, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
19BMSPhotometer TWC07/06/201219:35:0007/09/201210:18:00A GFCI on the North tracker post was tripped due to thunderstorms. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
20BMSPhotometer TWC05/26/201206:15:0005/26/201207:12:00Due to a planned power outage for SRRL electrical work, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, PGS-100, POM-01, TWC.
21BMSPhotometer TWC11/12/201111:59:0011/12/201113:26:00Due to an undervoltage followed by a power outage at 13:12, the following BMS instruments may be reading bogus or invalid data: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TWC.
22BMSPhotometer TWC09/08/201105:56:0009/08/201109:38:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, PGS-100, POM-01, TWC.
23BMSPhotometer TWC07/12/201122:57:0007/13/201107:40:00Due to a tripped GFCI outlet at SRRL the following instruments were without power during this time: TWC & POM-01 Photometers, Direct CUVA2 & CUVB2, and the PGS100. The readings are either zero or bogus and should not be used. GFCI was tripped due to Eppley venitlators which were water logged by thunderstorm.
24BMSPhotometer TWC06/14/201113:01:0006/14/201113:14:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
25BMSPhotometer TWC05/25/201105:19:0005/26/201112:16:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.